calculate the no of atoms of constituent element in 53g of NA2CO3

Grade:11

2 Answers

Sheenam
7 years ago
3.011×10^23 atoms  (use mole concept)
Atomic mass=avogardo no.
106g of Na2CO3=6.022*10^23atons
53g of Na2CO3=3.011*10^23atoms
khushi singh
6 years ago
Molar mass of Na2CO3 is106g.1 mole=6.022×10^23 particles106g of Na^2CO^3 contain=6.022×10^23molecules1g of Na^2CO^3contain = 6.022×10^23/10653 g of Na^2CO^3 contain=0.056×53×10^23moleculesAns=3.011×10^23molecules